OpenStudy (anonymous):

Solve by completing the square: -x^2+6x+10=0 Last one lol, please help.

OpenStudy (anonymous):

Help?lol, I know I have to add to both sides. I just need help lol.

OpenStudy (radar):

Multiply thru by -1 getting x^2 -6x-10=0. Look closely at the x^2 -6x what would be here x^2 -6x ________ to be a perfect square? How about x^2 -6x + 9 ???? that would be the same as (x-3)^2. To get a +9 I must add 19 which I will do:x^2 - 6x +9 = 19 After adding 19 to both sides. Follow so far??

OpenStudy (radar):

-10+19 = 9. Now we write it like this (x -3)^2 = 19 Take the square root of both sides getting \[x-3 = \sqrt{19}\]\[x = 3 \pm \sqrt{19}\]

OpenStudy (radar):

I hope you could follow that @ILoveLove

OpenStudy (anonymous):

Yes sir I do, for the final answer would be to the second equation you have?

OpenStudy (radar):

The final answer is\[x =3\pm \sqrt{19}\]Which would be spoken as 3 plus and minus the square root of 19